MrBibbers:

What is the axis of symmetry of h(x) = –2x2 + 12x – 3? x = –15 x = –3 x = 3 x = 15

Elsa213:

`the axis of symmetry is the x value of the vertex` `we have a handy-dandy way of finding that from standard form, ax^2+bx+c=y` `for` `ax^2+bx+c=y` `the x value of the vertex is -b/2a` `y=-2x^2+12x-3` `-b/2a=-12/(2*-2)=-12/-4=3` `x=3 is the axis of symmetry`

Elsa213:

Source: https://brainly.com/question/2038147

MrBibbers:

Nice copy and paste, milk jugs.

Elsa213:

I know. ;3

Ultrilliam:

you more and more appear to be historyguy by the second...